a bag contain a totap of 12 batteries of which six are defective. Selecting two at a random without replacement, determin the probabilty that none of the batteries you select are good.
Answer by stanbon(75887)

a bag contain a total of 12 batteries of which six are defective. Selecting two at a random without replacement, determine the probabilty that none of the batteries you select are good,
---
P(2 are not good) = 6C2/12C2 = 15/66 = 5/22
---------------------
Cheers,
Stan H.